# Tractable Equilibrium Computation in Markov Games through Risk Aversion

### v1

@@ 2.2 Bounded Rationality in Matrix Games
When all players achieve a Nash equilibrium in the space of their quantal responses— the class of strategies that can be represented by a fixed class of quantal response functions (cf. Definition 4)—the resulting equilibrium is known as a Quantal Response Equilibrium (QRE).
To incorporate quantal responses into our risk adjusted game we introduce regularization to the player’s losses. This can be shown to be equivalent to constraining the player’s responses to quantal responses (see e.g., [21, Proposition 7], or [69, 45]). The regularized risk adjusted game takes the form:
$\begin{split}f^{\epsilon_{i}}_{i}(\pi_{i},\pi_{-i})=\sup_{p_{i}\in\mathcal{P}_{-i}}-\pi_{i}^{T}R_{i}p_{i}-D_{i}(p_{i},\pi_{-i})+\epsilon_{i}\nu_{i}(\pi_{i})\\ \end{split}$ (6)
@@ 3 Conditions for Computational Tractability of RQE
To do so, we first introduce a related $2n$-player game in which we associate to each original player $i$ an adversary whose strategy we denote $p_{i}\in\mathcal{P}_{-i}$. Let $p=(p_{1},...,p_{n})\in\bar{\mathcal{P}}=\prod_{i=1}^{n}\mathcal{P}_{-i}$. In this new game, each original player’s loss function takes the form:
$J_{i}(\pi_{i},\pi_{-i},p)=-\pi_{i}^{T}R_{i}p_{i}-D_{i}(p_{i},\pi_{-i})+\epsilon_{i}\nu_{i}(\pi_{i})$
@@
Given the definition of the four player game we show how no-regret learning algorithms can be used to compute a RQE. To do so, we show that CCEs of the $2n$-player game coincide with NEs of it and NEs of the $2n$-player game coincide with RQE of the original $n$-player matrix game. This is a phenomenon known as equilibrium collapse which is well known to happen in zero-sum games and certain generalizations of zero-sum games [7, 38]. To prove this, we first define coarse correlated equilibria.
@@
$\mathbb{E}_{(\pi,p)\sim\sigma}[J_{i}(\pi,p)]$ $\leq\mathbb{E}_{(\pi_{-i},p)\sim\sigma}[J_{i}(\pi^{\prime}_{1},\pi_{-i},p)]\ \ \ \forall\pi^{\prime}_{i}\in\Delta_{A_{i}}$
@@
With this definition in hand we now present our results on computational tractability of RQE in 2-player matrix games. We also extend the results to $n$-player setting in Section B.1. To simplify our results we define $H_{1}(p_{1},\pi_{2})=D_{1}(p_{1},\pi_{2})-\xi_{1}\nu_{2}(\pi_{2})$ and $H_{2}(p_{2},\pi_{1})=D_{2}(p_{2},\pi_{1})-\xi_{2}\nu_{1}(\pi_{1})$. Let $\xi_{1}^{*}>0$ and $\xi_{2}^{*}>0$ be the smallest values of $\xi_{1}$ and $\xi_{2}$ such that $H_{1}(p_{1},\pi_{2})$ and $H_{2}(p_{2},\pi_{1})$ are concave in $\pi_{2}$ and $\pi_{1}$ for all $p_{1}$ and $p_{2}$ respectively. The parameters $\xi^{*}_{1}$ and $\xi^{*}_{2}$ capture the player’s degrees of risk aversion.
The following theorem gives conditions on $\xi^{*}_{1},\xi^{*}_{2}$ and $\epsilon_{1},\epsilon_{2}$, the bounded rationality parameters under which RQE’s are computationally tractable using no-regret learning. The proof is postponed to Appendix C.
@@
The previous theorem gives a range of risk aversion and bounded rationality parameters under which an RQE can be computationally tractable using no-regret learning. Before presenting corrolaries for certain instantiations of risk measures and quantal response functions in Appendix B, we briefly remark that the result does not necessarily guarantee uniqueness, though by exploiting the connections between socially convex games and monotone games [25] such a result would follow.
To validate this result, we show that the class of RQE that is computationally tractable seems sufficiently rich to capture the aggregate strategy played by people in various matrix games studied in laboratory experiments in behavioral economics [60, 27, 61] to showcase differing human behavior from Nash equilibrium strategies. In particular, we showcase in Fig. 1 that for some choice of parameters ($\epsilon_{j}$ and $\tau_{j}$’s) satisfying conditions in Theorem 3, we recover risk-averse and rational human behaviors in various matching pennies games mentioned in Tables 3 and 2. We provide more details about the games and experiments in Section E.1 for reproducing our results.
@@ 4.2 Algorithm and Theoretical Guarantees
The proof is postponed to Appendix D. We remark that $L$ can be some constant for various penalty function, such as $L=1$ when $\{D_{\mathsf{env},i}\}_{i\in[n]}$ are defined as any $\ell_{p}$ norm including total variation (TV). We also remark that similar sample complexity guarantees can be shown when $\{D_{\mathsf{env},i}\}_{i\in[n]}$ are defined as $\phi$-divergence and allude to analyses ideas in [51, 75, 66]. Note that computing exact RQE can be computationally expensive or may not be necessary in practice. This result provides the first finite-sample guarantees of computing RQE in RAMGs. We show that even when the environment is not perfectly known but needs to be learned by samples, we can achieve approximate-RQEs of RAMGs using sufficient finite samples of size $N_{\mathsf{all}}$.
@@ 4.3 Experiments and Evaluation
We consider a grid-world problem to test our algorithm (Algorithm 1) for showcasing risk-averse and bounded rational behaviors. We note that we use our matrix game method (details described in Section E.1) for the $\mathsf{RQE}$ solver at Line 8 in Algorithm 1.
@@
We use the penalty function KL with the log-barrier regularizer, as mentioned in Tables 1 and 2, for our experiments. We also use KL divergence for the environmental uncertainty metric $D_{\mathsf{env},i}$ with the same bounded rational parameter $\tau_{i}^{-1}$ used by $\mathsf{RQE}$. We use the full information of this grid-world to evaluate the Q-values described in Algorithm 1. Additionally, as we use the matrix games solver for $\mathsf{RQE}$, we have similar statistical deviations for different training runs. In Section E.2, we also showcase results with $\ell_{1}$ metric for the environmental uncertainty metric $D_{\mathsf{env},i}$.
@@ Acknowledgment
Appendix A Relating Nash of the $2n$-player game and RQE
We first note that this $2n$ player game is a convex game and that a Nash equilibrium exists. To relate outcomes in the new game to that of the original game, we show that the strategies played by the original players in Nash equilibria of this new game coincide with Nash equilibria of the original risk-adjusted game—i.e., the RQE.
@@ Appendix A Relating Nash of the $2n$-player game and RQE
Appendix B Further results on computational tractability of RQE in Matrix Games
To begin, we present two corollaries that specialize the results to specific risk metrics and quantal responses. In the first we look at the case where players make use of the entropic risk and log-barrier reguarlizers.
Corollary 5.1.
Corollary 5.2.
@@ Appendix B Further results on computational tractability of RQE in Matrix Games
B.1 Computing RQE in n-player General-Sum Games
To do so we now define $H_{i}(p_{i},\pi_{-i})=D_{i}(p_{i},\pi_{-i})-\sum_{j\neq i}\xi_{i,j}\nu_{j}(\pi_{j})$.
For all $i,j\in\{1,...,n\}$ let $\xi_{i,j}^{*}>0$ be the smallest values of $\xi_{i,j}$ such that $H_{i}(p_{i},\pi_{-i})$ is concave in $\pi_{j}$. Again, the parameters $\xi^{*}_{i,j}$ capture the player’s degrees of risk aversion. The following theorem gives a general condition under which an RQE is computable using no-regret learning.
Theorem 6.
Proof.
We can now take a sum of the $2n$ utility functions to find that:
$=\sum_{i=1}^{n}J_{i}(\hat{\pi},\hat{p})+\bar{J}_{i}(\hat{\pi},\hat{p})$
Thus we have shown that:
This implies that:
Thus we have shown that $(\hat{\pi},\hat{p})$ is a Nash equilibrium for the $2n$-player game and that consequently $\hat{\pi}$ must be a RQE for the risk-adjusted game. ∎
Appendix C Proof of Theorem 3
@@
We note a fact: $\lambda a+(1-\lambda)b=\lambda c+(1-\lambda)d$ and $a\leq c,b\leq d$ implies $a=c,b=d$ for any $a,b,c,d\in\mathbb{R}$. Using this fact for Eqs. 23 and 24, we have:
@@
Thus we have shown that $\hat{z}=(\hat{\pi}_{1},\hat{p}_{1},\hat{\pi}_{2},\hat{p}_{2})$ is a Nash equilibrium for the 4-player game and that consequently $(\hat{\pi}_{1},\hat{\pi}_{2})$ must be a RQE for the risk-adjusted game. $\square$
Appendix D Proof of Theorem 5

### v2

@@ 2.2 Bounded Rationality in Matrix Games
When all players achieve a Nash equilibrium in the space of their quantal responses—the class of strategies that can be represented by a fixed class of quantal response functions (cf. Definition 4)—the resulting equilibrium is known as a Quantal Response Equilibrium (QRE). To incorporate quantal responses into our risk-adjusted game, we introduce regularization to the player’s losses. This can be shown to be equivalent to constraining the player’s responses to quantal responses (see, e.g., [23, Proposition 7], or [68, 47]). We write a regularized risk-adjusted game as:
$\begin{split}f^{\epsilon_{i}}_{i}(\pi_{i},\pi_{-i})=f_{i}(\pi_{i},\pi_{-i})+\epsilon_{i}\nu_{i}(\pi_{i})\\ \end{split}$ (8)
@@
For any set of convex regularizers $\nu_{i}$, it is easy to observe the that game remains a convex game and thus RQE will exist.
@@ 3 Conditions for Computational Tractability of RQE
Remark 1.
For ease of exposition we focus on deriving the results for the aggregate risk-aversion formulation and defer the results for action-dependent risk-aversion to Appendix B. The results for both formulations in terms of computational tractability are similar in that both allow for a class of RQE to be learned using no-regret learning. However, we remark that the analysis of the aggregate case is comparatively simpler and allows one to compute simple bounds on the set of RQE that are computationally tractable. The action-dependent case requires more work and slightly more stringent requirements on the relationship between bounded rationality and risk-aversion.
To derive our results on the computational tractability of RQE under aggregate risk-aversion, we first introduce a related $2n$-player game in which we associate to each original player $i$ an adversary whose strategy we denote $p_{i}\in\mathcal{P}_{-i}$. Let $p=(p_{1},...,p_{n})\in\bar{\mathcal{P}}=\prod_{i=1}^{n}\mathcal{P}_{-i}$. In this new game, each original player’s loss function takes the form:
$J_{i}(\pi_{i},\pi_{-i},p)=-\pi_{i}^{T}R_{i}p_{i}-D_{i}(p_{i},\pi_{-i})+\epsilon_{i}\nu_{i}(\pi_{i})$ (9)
@@
Given the definition of the $2n$-player game we show how no-regret learning algorithms can be used to compute a RQE. To do so, we show that CCEs of the $2n$-player game coincide with NEs of it and NEs of the $2n$-player game coincide with RQE of the original $n$-player matrix game. This is a phenomenon known as equilibrium collapse which is well known to happen in zero-sum games and certain generalizations of zero-sum games [10, 38]. To prove this, we first define coarse correlated equilibria.
@@
$\mathbb{E}_{(\pi,p)\sim\sigma}[J_{i}(\pi,p)]$ $\leq\mathbb{E}_{(\pi_{-i},p)\sim\sigma}[J_{i}(\pi^{\prime}_{i},\pi_{-i},p)]\ \ \ \forall\pi^{\prime}_{i}\in\Delta_{A_{i}}$
@@
With this definition in hand we now present our results on computational tractability of RQE in 2-player matrix games. We also extend the results to $n$-player setting in Section A.3.
To simplify our results we define $H_{1}(p_{1},\pi_{2})=D_{1}(p_{1},\pi_{2})-\xi_{1}\nu_{2}(\pi_{2})$ and $H_{2}(p_{2},\pi_{1})=D_{2}(p_{2},\pi_{1})-\xi_{2}\nu_{1}(\pi_{1})$. Let $\xi_{1}^{*}>0$ and $\xi_{2}^{*}>0$ be the smallest values of $\xi_{1}$ and $\xi_{2}$ such that $H_{1}(p_{1},\pi_{2})$ and $H_{2}(p_{2},\pi_{1})$ are concave in $\pi_{2}$ and $\pi_{1}$ for all $p_{1}$ and $p_{2}$ respectively. The parameters $\xi^{*}_{1}$ and $\xi^{*}_{2}$ capture the player’s degrees of risk aversion as illustrated in Corollaries 6.1 and 6.2 in Appendix A respectively.
The following theorem gives conditions on $\xi^{*}_{1},\xi^{*}_{2}$ and $\epsilon_{1},\epsilon_{2}$, the bounded rationality parameters under which RQE’s are computationally tractable using no-regret learning. The proof is postponed to Section A.2.
@@
Assume the penalty functions that give rise to the players’ risk preferences $D_{1}(\cdot,\cdot)$ and $D_{2}(\cdot,\cdot)$ are jointly convex in both their arguments. If $\sigma$ is a CCE of the four player game with $\xi_{1,2}=\xi_{1}^{*}$ and $\xi_{2,1}=\xi_{2}^{*}$, and
@@
The previous theorem gives a range of risk aversion and bounded rationality parameters under which an RQE can be computationally tractable using no-regret learning. Before presenting corollaries for certain instantiations of risk measures and quantal response functions in Appendix A, we briefly remark that the result does not necessarily guarantee uniqueness, though by exploiting the connections between socially convex games and monotone games [26] such a result would follow.
To validate this result, we show that the class of RQE that is computationally tractable seems sufficiently rich to capture the aggregate strategy played by people in various matrix games studied in laboratory experiments in behavioral economics [58, 28, 59] to showcase differing human behavior from Nash equilibrium strategies. In particular, we showcase in Fig. 1 that for some choice of parameters ($\epsilon_{j}$ and $\tau_{j}$’s) satisfying conditions in Theorem 3, we recover risk-averse and rational human behaviors in various matching pennies games mentioned in Tables 3 and 2. We provide more details about the games and experiments in Section D.1 for reproducing our results.
@@ 4.2 Algorithm and Theoretical Guarantees
The proof is postponed to Appendix C. We remark that $L$ can be some constant for various penalty function, such as $L=1$ when $\{D_{\mathsf{env},i}\}_{i\in[n]}$ are defined as any $\ell_{p}$ norm including total variation (TV). We also remark that similar sample complexity guarantees can be shown when $\{D_{\mathsf{env},i}\}_{i\in[n]}$ are defined as $\phi$-divergence and allude to analyses ideas in [54, 73, 64, 65]. Note that computing exact RQE can be computationally expensive or may not be necessary in practice. This result provides the first finite-sample guarantees of computing RQE in RAMGs. We show that even when the environment is not perfectly known but needs to be learned by samples, we can achieve approximate-RQEs of RAMGs using sufficient finite samples of size $N_{\mathsf{all}}$.
@@ 4.3 Experiments and Evaluation
We consider a grid-world problem to test our algorithm (Algorithm 1) for showcasing risk-averse and bounded rational behaviors. We note that we use our matrix game method (details described in Section D.1) for the $\mathsf{RQE}$ solver at Line 8 in Algorithm 1.
@@
We use the penalty function KL with the log-barrier regularizer, as mentioned in Tables 1 and 2, for our experiments. We also use KL divergence for the environmental uncertainty metric $D_{\mathsf{env},i}$ with the same bounded rational parameter $\tau_{i}^{-1}$ used by $\mathsf{RQE}$. We use the full information of this grid-world to evaluate the Q-values described in Algorithm 1. Additionally, as we use the matrix games solver for $\mathsf{RQE}$, we have similar statistical deviations for different training runs. In Section D.2, we also showcase results with $\ell_{1}$ metric for the environmental uncertainty metric $D_{\mathsf{env},i}$.
@@ Acknowledgment
Appendix A Further results on computational tractability of RQE in aggregate risk-averse matrix games
In this section we provide further results on the aggregate risk aversion formulation of the risk-averse problem. This section provides a proof of the main result of Theorem 3 presented in the main body of the paper as well as the generalization of the results on computationaly tractability to $n$-players. A key step in the proof is relating the Nash equilibrium of the $2n$-player convex game presented where players have utilites of the form (10) and (9) and RQE of the desired game (3). We then use this result to prove Theorem 3.
A.1 Relating Nash of the $2n$-player game and RQE
We first note that this $2n$ player game is a convex game played over compact convex action sets and thus a Nash equilibrium must exist [56]. To relate outcomes in this new game to that of the original game, we show that the strategies played by the original players in Nash equilibria of the $2n$-player game coincide with the RQE.
@@ A.1 Relating Nash of the $2n$-player game and RQE
A.2 Proof of Theorem 3 and immediate corollaries
Given the results showing that finding a Nash equilibrium of the $2n$-player convex game allows us to compute RQE, we present the proof of Theorem 3. We also provide two corollaries that specialize the results to different risk-metrics and quantal response functions.
Theorem 6 (Restatement of Theorem 3).
$\frac{\epsilon_{1}}{\xi^{*}_{1}}\geq\frac{\xi_{2}^{*}}{\epsilon_{2}},$
The proof follows by showing that CCE of the $2n$-player convex game coincide with Nash equilibria of the $2n$-player convex game, and then invoking results from the previous section.
@@ A.2 Proof of Theorem 3 and immediate corollaries
We note a fact: $\lambda a+(1-\lambda)b=\lambda c+(1-\lambda)d$ and $a\leq c,b\leq d$ implies $a=c,b=d$ for any $a,b,c,d\in\mathbb{R}$. Using this fact for Eqs. 26 and 27, we have:
@@
Thus we have shown that $\hat{z}=(\hat{\pi}_{1},\hat{p}_{1},\hat{\pi}_{2},\hat{p}_{2})$ is a Nash equilibrium for the 4-player game. By invoking Propostion 1 we derive our result that $(\hat{\pi}_{1},\hat{\pi}_{2})$ must be a RQE for the original risk-adjusted game. ∎
We now present two corollaries that specialize the results to specific risk metrics and quantal responses. In the first we look at the case where players make use of the entropic risk and log-barrier reguarlizers.
Corollary 6.1.
Corollary 6.2.
Proof.
A.3 Computing RQE in n-player General-Sum Games
To do so we now define $H_{i}(p_{i},\pi_{-i})=D_{i}(p_{i},\pi_{-i})-\sum_{j\neq i}\xi_{i,j}\nu_{j}(\pi_{j})$. For all $i,j\in\{1,...,n\}$ let $\xi_{i,j}^{*}>0$ be the smallest values of $\xi_{i,j}$ such that $H_{i}(p_{i},\pi_{-i})$ is concave in $\pi_{j}$. Again, the parameters $\xi^{*}_{i,j}$ capture the player’s degrees of risk aversion. The following theorem gives a general condition under which an RQE is computable using no-regret learning.
Theorem 7.
Proof.
We can now take a sum of the $2n$ utility functions to find that:
$=\sum_{i=1}^{n}J_{i}(\hat{\pi},\hat{p})+\bar{J}_{i}(\hat{\pi},\hat{p})$
Thus we have shown that:
This implies that:
Thus we have shown that $(\hat{\pi},\hat{p})$ is a Nash equilibrium for the $2n$-player game. By invoking Proposition 1 we can observe that $\hat{\pi}$ must be a RQE for the risk-adjusted game. ∎
Appendix B Computational tractability of RQE in action-dependent risk-averse matrix games
As in the case of aggregate risk-aversion, to prove our results we again introduce an auxiliary game that we relate to our risk-adjusted game of interest. In this case, the loss of the original players is given by:
$J_{i}(\pi_{i},\pi_{-i},p)=\sum_{j\in\mathcal{A}_{i}}\pi_{i}(j)\left(-\langle R_{i,j},p_{i,j}\rangle-D_{i}(p_{i,j},\pi_{-i})\right)+\epsilon_{i}\nu_{i}(\pi_{i}).$ (28)
We now associate each player $i$ to its intermediate adversary $p_{i}$. For each player $p_{i}$ their loss function is given by:
$\bar{J}_{i}(\pi,p_{i})$ $=\sum_{j\in\mathcal{A}_{i}}\pi_{i}(j)\left(\langle R_{i,j},p_{i,j}\rangle+D_{i}(p_{i,j},\pi_{-i})\right)-\sum_{k}\xi_{i,k}\nu_{k}(\pi_{k}),$ (29)
where $p_{i}=\{p_{i,j}\}_{j\in\mathcal{A}_{i}}$ where $p_{i,j}\in\mathcal{P}_{-i}$. This is once again a convex game since each player’s loss is convex in its own argument. Define $\xi_{i,k}^{*}\geq 0$ as the minimum value of $\xi_{i,k}$ needed for $\bar{J}_{i}(\pi,p_{i})$ to be concave in $\pi$ for all values of $p_{i}$ for all $i$. Note that due to the structure of $\bar{J}$ the values of $\xi_{i,k}^{*}$ only depend on properties of the risk metrics under consideration which are captured in $D_{i}$, and not on the payoff structure $R_{i}$.
The following proposition relates the Nash equilibrium of the $2n$-player convex game to the RQE of the action-dependent risk averse game.
Proposition 2.
If $(\pi^{*},p^{*})$ is a Nash equilibrium of the $2n$-player game, then $\pi^{*}$ is a RQE of Game (8). Furthermore, if $\pi^{*}$ is a RQE of the action-dependent risk averse game, then $(\pi^{*},p^{*})$ is a Nash equilibrium of the $2n$-player game, where:
$p_{i,j}^{*}=\arg\min_{p_{i,j}\in\mathcal{P}_{-i}}\bar{J}_{i,j}(\pi^{*},p_{i,j})$ (30)
The proof of this result follows by exactly the same arguments as the proof of Proposition 1 and is therefore omitted for brevity.
Given these results we now prove the equivalent of Theorem 7 for the action-dependent formulation. We note that by a more careful accounting of terms, a stronger guarantee is possible in the $2$-player regime similar to Theorem 3.
Theorem 8.
Assume the penalty functions that give rise to the players’ risk preferences $D_{i}(\cdot,\cdot)$ are jointly convex in both their arguments. If $\sigma$ is a CCE of the $n$-player game and for each $i$:
$\epsilon_{i}\geq\sum_{j}\xi_{i,j}^{*}$
then the marginal strategies $\hat{\pi}_{i}=\mathbb{E}_{\sigma}[\pi_{i}]$ constitute a RQE of the action-dependent risk-averse game.
Note that one point of departure for this result from the previous results is that $\sigma$ is now the CCE of the original 2-player convex game defined on the objective functions $f_{1}^{\epsilon_{1}}$ and $f_{2}^{\epsilon_{2}}$.
Proof.
To begin, let $\sigma$ be a CCE of the convex game played on $f_{i}^{\epsilon_{i}}$, where:
$f^{\epsilon_{i}}_{i}(\pi_{i},\pi_{-i})=\sum_{j\in\mathcal{A}_{i}}\pi_{i}(j)\left(\sup_{p_{i,j}\in\mathcal{P}_{-i}}-\langle R_{i,j},p_{i,j}\rangle-D_{i}(p_{i,j},\pi_{-i})\right)+\epsilon_{i}\nu(\pi_{i}).$
By definition of a CCE, we must have that:
$\mathbb{E}_{\sigma}[f^{\epsilon_{i}}_{i}(\pi_{i},\pi_{-i})]\leq\mathbb{E}_{\sigma}[f^{\epsilon_{i}}_{i}(\pi_{i}^{\prime},\pi_{-i})]\ \ \forall\ \pi_{i}^{\prime}\in\Delta_{A_{i}},$
for all $i=1,...,n$. Given $\sigma$, we can define a new distribution $\sigma^{\prime}$ as $(\pi,p^{*}(\pi))$ where
$p_{i}^{*}(\pi)=\arg\min_{p_{i}\in\mathcal{P}_{-i}}\bar{J}_{i}(\pi,p_{i}),$
with $\xi_{i,j}=\xi_{i,j}^{*}$. We now claim that, by construction, $\sigma^{\prime}$ is a CCE of the $2n$-player game. To see this, we observe that
$\mathbb{E}_{\sigma^{\prime}}[J_{i}(\pi_{i},\pi_{-i},p)]=\mathbb{E}_{\sigma}[f^{\epsilon_{i}}_{i}(\pi_{i},\pi_{-i})]$ $\leq\mathbb{E}_{\sigma}[f^{\epsilon_{i}}_{i}(\pi_{i}^{\prime},\pi_{-i})]=\mathbb{E}_{\sigma^{\prime}}[J_{i}(\pi_{i}^{\prime},\pi_{-i},p)]\ \ \forall\ \pi_{i}\in\Delta_{A_{i}},$
where the first equality follows by construction of $\sigma^{\prime}$, the second from the definition of a CCE, and the third from the fact that the value of $p^{*}(\pi)$ only depends on $\pi_{-i}$ and not $\pi_{i}^{\prime}$.
Similarly, we can show the same result for all the players $p_{i}$. By simply applying Jensen’s inequality, we can see that
$\mathbb{E}_{\sigma^{\prime}}[\bar{J}_{i}(\pi,p_{i})]=\mathbb{E}_{\pi\sim\sigma}\left[\min_{p_{i,j}}\bar{J}_{i}(\pi,p_{i,j})\right]\leq\mathbb{E}_{\sigma^{\prime}}[\bar{J}_{i}(\pi,p^{\prime}_{i})]\ \ \forall p^{\prime}_{i}\in\mathcal{P}_{-i}.$
Thus we can observe that $\sigma^{\prime}$ is a CCE of the $2n$-player game. To show that the marginals of $\pi$ in this CCE are Nash equilibria of the convex game (and thus RQE), we proceed as before. Using the fact that $D(p_{i,j},\pi_{-i})$ is jointly convex in each of its arguments, we can apply Jensen’s inequality to find that:
$\mathbb{E}_{\sigma^{\prime}}[J_{i}(\pi_{i},\pi_{-i},p)]\leq J_{i}(\pi_{i}^{\prime},\hat{\pi}_{-i},\hat{p})\ \ \ \forall\pi_{i}^{\prime}\in\Delta_{A_{i}},$
where $\hat{p}_{i,j}=\mathbb{E}_{\sigma^{\prime}}[p_{i,j}]$ and $\hat{\pi}_{i}=\mathbb{E}_{\sigma^{\prime}}[\pi_{i}]$. Similarly, by our choice of $\xi_{i,j}=\xi_{i,j}^{*}$ we can find that:
$\mathbb{E}_{\sigma^{\prime}}[\bar{J}_{i}(\pi,p_{i})]\leq\bar{J}_{i}(\hat{\pi},p_{i}^{\prime})\ \ \ \forall p_{i}^{\prime}\in\mathcal{P}_{-i}.$
Now, we can observe that:
$\sum_{i}J_{i}(\hat{\pi}_{i},\hat{\pi}_{-i},\hat{p})+\bar{J}_{i}(\hat{\pi},\hat{p}_{i})$ $\geq\mathbb{E}_{\sigma^{\prime}}\left[\sum_{i}J_{i}(\pi_{i},\pi_{-i},p)+\bar{J}_{i}(\pi,p_{i})\right]$
$=\sum_{i}\mathbb{E}_{\sigma^{\prime}}\left[\left(\epsilon_{i}-\sum_{j}\xi_{i,j}^{*}\right)\nu_{i}(\pi_{i})\right]$
$\geq\sum_{i}\left(\epsilon_{i}-\sum_{j}\xi_{i,j}^{*}\right)\nu_{i}(\hat{\pi}_{i})$
$=\sum_{i}J_{i}(\hat{\pi}_{i},\hat{\pi}_{-i},\hat{p})+\bar{J}_{i}(\hat{\pi},\hat{p}_{i}),$
where the third inequality follows by the Jensen’s inequality. By the same rationale as in the proof of Theorem 7 we can now conclude that $\hat{\pi}$ is a Nash equilibrium-joint strategy profile-of the $2n$ player game and thus (due to Proposition 2) an RQE of the action-dependent risk averse game. ∎
Similar to the aggregate risk regime, one can instantiate the previous theorem with specific risk-metrics and quantal response functions to illustrate the class of action-dependent RQE that are guaranteed to be computationally tractable. Importantly, this class once again only depends on the levels of risk averse and quantal response but not on the underlying structure of the game. However, note that the requirements are more stringent since the requirements on the $\xi^{*}$’s are stronger since they must ensure joint convexity of $\bar{J}_{i}$. Nevertheless further algebraic manipulations would allow one to recover analogues of Corollaries 6.1 and 6.2 for the action-dependent risk case as well. For brevity we leave these as exercises to the reader.
Appendix C Proof of Theorem 5